### v3

# FedCC: Robust Federated Learning against Model Poisoning Attacks

Abstract
Federated learning is a distributed framework designed to address privacy concerns. However, it introduces new attack surfaces, which are especially prone when data is non-Independently and Identically Distributed. Existing approaches fail to effectively mitigate the malicious influence in this setting; previous approaches often tackle non-IID data and poisoning attacks separately. To address both challenges simultaneously, we present FedCC, a simple yet effective novel defense algorithm against model poisoning attacks. It leverages the Centered Kernel Alignment similarity of Penultimate Layer Representations for clustering, allowing the identification and filtration of malicious clients, even in non-IID data settings. The penultimate layer representations are meaningful since the later layers are more sensitive to local data distributions, which allows better detection of malicious clients. The sophisticated utilization of layer-wise Centered Kernel Alignment similarity allows attack mitigation while leveraging useful knowledge obtained. Our extensive experiments demonstrate the effectiveness of FedCC in mitigating both untargeted model poisoning and targeted backdoor attacks. Compared to existing outlier detection-based and first-order statistics-based methods, FedCC consistently reduces attack confidence to zero. Specifically, it significantly minimizes the average degradation of global performance by 65.5%. We believe that this new perspective on aggregation makes it a valuable contribution to the field of FL model security and privacy. The code will be made available upon acceptance.
Keywords:
Federated learning model poisoning attack backdoor attack robust aggregation
1 Introduction
Federated Learning (FL) [22] is a distributed model training framework designed to preserve privacy. FL allows data to remain on client devices but exchanges their local model parameters only for model training. It thus reduces (data) privacy risks associated with centralized learning, not requiring central data collection. Thus, FL is especially suitable for data-sensitive environments since raw data remains on clients, reducing the risk of data leakage on the server or in between communication.
However, its distributed nature makes FL susceptible to model poisoning attacks [12]. The server cannot directly examine local datasets’ data quality or model parameter integrity, leaving compromised clients or attackers to manipulate local models. This can degrade global model performance indiscriminately (untargeted attacks) [7, 2, 26] or cause incorrect predictions on specific inputs (targeted attacks) [29, 31, 3, 33]. Backdoor attacks are stealthier targeted attacks that maintain overall performance while misclassifying specific inputs [1].
Some preliminary defenses propose robust aggregation methods, but they may compromise privacy by sharing raw data or exposing data distribution to untrusted parties [6, 30]. While some defenses exist, most focus on IID data settings, which are rare in real-world scenarios. Non-IID data, with variations in class distribution and features across clients [12], can gradually degrade model performance and complicate the detection of malicious clients [39, 35], as the impact of attacks increases with the degree of non-IID [19, 26].
In this study, we introduce FedCC, a defense mechanism that mitigates untargeted model poisoning and targeted backdoor attacks in both IID and non-IID settings. FedCC leverages the Centered Kernel Alignment (CKA) of Penultimate Layer Representations (PLRs) to distinguish malicious clients from benign ones. PLRs are known to be highly specific to local data distributions [36], making clients distinguishable, especially when data is non-IID. Exploiting CKA’s well-known capability of measuring similarity in high-dimensional space, such as neural network representations, we accurately measure the similarity between the PLR of different clients, further allowing more accurate detection of malicious clients. Empirical evidence supporting our rationale is shown in Figure 2 and Table 1. The use of PLRs shows the highest separability between benign and malicious clients in all settings, as seen by the largest distance between clusters in the PLR. The use of CKA, specifically kernel CKA, is justified by its highest performance in both targeted and untargeted IID and non-IID settings. Furthermore, our method maintains data privacy by using similarity measures between local and global models, not having shared data. We named our proposed method FedCC, Federated learning using Clustered CKA.
Empirical experiments show that FedCC outperforms existing defenses that rely on first-order statistics, which usually suffer from a high false negative detection rate. It significantly improves global accuracy against untargeted attacks (66.27% compared to similar approaches) and reduces attack confidence in targeted backdoor attacks to almost zero while preserving main task accuracy. These results highlight FedCC’s effectiveness and robustness in mitigating attacks in FL settings, particularly in non-IID data settings.
Our contributions are summarized as follows:
  • We propose FedCC, a novel and scalable defense method against model poisoning attacks using the CKA of PLRs of local and global models and performing layer-wise weighted aggregation of their model parameters.
  • We justify that PLR is a highly distinguishable feature to detect malicious clients by measuring the discrepancy between two clusters of clients.
  • We demonstrate that CKA provides the most sensitive and accurate similarity measure between differently trained models when the server has no knowledge about clients or centralized data, even in non-IID settings.
  • We empirically demonstrate the effectiveness of FedCC through exhaustive experiments. Our results show that comparing the CKA of PLRs filters malicious clients more reliably than the first-order statistics. FedCC outperforms existing defenses even in non-IID settings.
2 Backgrounds and Related Works
2.1 Poisoning Attacks in FL
FL is vulnerable to poisoning attacks due to its distributed nature, where the server cannot directly inspect dataset quality or model integrity. Model poisoning attacks are particularly destructive, as adversaries can stealthily manipulate local model parameters to degrade global model performance [3, 5], and this paper focuses on such attacks.
In model poisoning attacks, client-side adversaries alter local model parameters before submitting them to the server. To avoid detection and prevent global model divergence, attackers optimize the local models for both training loss and an adversarial objective [5, 7]. For instance, A3FL [37] dynamically fine-tunes backdoor triggers to make them harder to detect, while IBA [24] generates robust triggers using a generative network, exploiting the global model as a discriminator. Attackers can further manipulate hyperparameters, such as learning rate, local epochs, batch size, and regularization [17, 21], dynamically before and during local training to evade detection. In a similar vein, DBA [33] embedded split triggers into local training data, associating them with targeted incorrect predictions. Unlike centralized backdoor attacks, DBAs distribute malicious updates across multiple adversarial participants, making them harder to detect with anomaly detection techniques.
Model poisoning attacks can be classified into untargeted and targeted types. Untargeted attacks degrade global model accuracy, while targeted attacks mislead the model to misclassify specific inputs without affecting other classes. Our defense mechanism mitigates both untargeted attacks by restoring global model accuracy and targeted backdoor attacks by overcoming the stealthiness and detection difficulty [31, 29].
2.2 Robust Aggregation algorithm in FL
Several Byzantine-robust aggregation methods, based on summary statistics or anomaly detection, have been proposed to mitigate model poisoning and backdoor attacks in FL. Krum [4], for instance, selects the update with the smallest Euclidean distance to others but assumes IID data and overlooks outliers [2]. Multi-Krum, which averages updates from multiple clients, is known to be more effective in non-IID settings [4]. Similarly, Median [34] computes the coordinate-wise median, Trimmed Mean [34] excludes extreme values, and Bulyan [10] combines Krum and Trimmed Mean for added robustness. These methods, however, require knowledge of the number of attackers, often unavailable in practice.
Inspired by the observation that malicious clients’ model parameters exhibit higher similarity, Foolsgold [8] identified Sybils by measuring cosine similarity between gradients. While effective against multiple Sybils, it struggled when only a single malicious client exists or with IID data, where it can overfit. FLTrust [6] addressed this by assigning trust scores based on ReLU-clipped cosine similarities but risking privacy by raw data sharing. Similarly, Lockdown [11] improved robustness by pruning parameters unused by the majority of clients during training, while FLIP [38] rejected low-confidence samples at test time in addition to the adversarial training.
Despite these advances, many methods rely on summary statistics (mean or median), direction (cosine similarity), or distance (Euclidean) of weight vectors. However, they often struggle to differentiate benign clients with non-IID data from malicious ones, leading to misclassification and suboptimal performance. The limited performance might be partially attributed to the inadequacy of Euclidean geometry for comparing neural network representations.
In contrast, our approach uses Kernel CKA, a kernel-based metric, to measure the similarity between global and local models. This method effectively identifies compromised model parameters among clients with diverse data distributions, enhancing defense against model poisoning in FL and overcoming the shortcomings of previous techniques.
2.3 Penultimate Layer Representation in FL
The penultimate layer is a neural network model’s second-to-last layer (i.e., before the softmax layer). Wang et al. [30] discovered that PLRs are effective in distinguishing malicious models from benign ones; within benign clients, benign PLRs follow the same distribution, whereas malicious PLRs do not, across datasets and neural network architectures. Specifically, they demonstrated that distances among benign PLRs are smaller than those of benign and malicious PLRs. Their proposed method, FLARE, assigned a trust score to each client based on pairwise PLR discrepancies defined by Maximum Mean Discrepancy among all model updates, allocating lower values to those farther from the benign distribution. The model updates are then scaled and averaged, weighted by each client’s trust score. Meanwhile, FLARE continuously redirects the global model using server-owned raw data, increasing the risk of data leakage and jeopardizing data privacy.
2.4 Centered Kernel Alignment
Centered Kernel Alignment (CKA) [13] is a highly accurate similarity metric that measures how similar two differently initialized or trained neural networks are, producing a value between 0 (not similar) and 1 (identical). CKA can be calculated by: $CKA(X,Y)=\frac{HSIC(X,Y)}{\sqrt{HSIC(X,X)HSIC(Y,Y)}}=\frac{{\lVert K(X^{T})K(Y)\rVert}^{2}}{{\lVert K(X^{T})K(X)\rVert}{\lVert K(Y^{T})K(Y)\rVert}}$ where $HSIC$ is Hilbert-Schmidt Independence Criterion [9] (i.e., a not normalized version of CKA) and $K$ is the RBF kernel. It constructs similarity kernel matrices in the weight space and compares these matrices to characterize the representation space, allowing comparisons between layers with varying widths and different initialization schemes.
CKA includes several vital properties for similarity measures of neural networks. It is non-invariant to invertible linear transformations, meaning similarity scores change if such transformations are applied. This property is vital; the gradient descent algorithm is not invariant to these transformations because similarity metrics invariant to linear transformations are inaccurate on models trained with gradient descent. It is also invariant to orthogonal transformations and isotropic scaling, which are desirable for neural networks trained with gradient descent owing to their stochastic nature. These properties make CKA a precise and reliable metric for measuring similarities between neural networks, particularly when comparing potential models from malicious clients.
CKA has been utilized to address data heterogeneity issues, taking into account that the similarity of non-IID data is notably lower than that of IID data [28], particularly in specific layers [27]. However, CKA has yet to be leveraged as a defense mechanism against attacks involving models from malicious clients where the data are distributed across the clients in a non-IID manner.
3 Threat Model
3.1 Attackers’ Goals
We consider two primary attack strategies in the context of FL: untargeted model poisoning and targeted backdoor attacks. Untargeted attacks (Fang attacks [7]) aimed to evade robust aggregation rules like Krum and Coordinate-wise Median or Trimmed Mean. Attackers manipulate the model parameters to degrade the global model’s overall accuracy. Targeted backdoor attacks, represented by Bhagoji et al. [3] and DBA [33], aim to deceive the global model into misclassifying specific data samples the attacker chose. The objective is to assign a different label chosen by the attacker while maintaining high accuracy for the remaining classes, making the attack more inconspicuous.
3.2 Attackers’ Capability
  • An attacker on the client side can control multiple compromised clients.
  • An attacker has full control over at most $k<n/2$ clients.
  • An attacker has knowledge about compromised clients’ data, such as current local model, previous global model, and hyperparameters.
  • An attacker has no knowledge or control over the server and honest clients.
3.3 Attack Strategy
3.3.1 Untargeted Model Poisoning Attacks [7]
specifically aim to break Krum and Coordinate-wise Median (Coomed) aggregation rules, known to be byzantine failure tolerant. Their primary objective is to disrupt the model training process, thereby diminishing global test accuracy. The attacks involve manipulating model parameters, such as flipping the sign of malicious parameters, to steer the model in the opposite direction from its uninterrupted trajectory. Specifically, we denote the attack against Krum as Untargeted-Krum and the attack against Coomed as Untargeted-Med.
Untargeted-Krum [7] alters malicious parameters resembling benign ones to maximize the chances of being selected by Krum. Specifically, the optimization problem is to find the maximum $\lambda,s.t.$ $w_{1}=Krum(w_{1},...,w_{m},w_{m+1},...,w_{n}),w_{1}=w_{G}-\lambda s,w_{i}=w_{1},$ for $i=2,3,...,m$ where $m$ is the number of attackers, $n$ is the total number of selected clients, $w_{G}$ is the previous global model, and $s$ is the sign of the global model parameter with no attack. The upper bound of the $\lambda$ can be solved as follows: $\lambda\leq\frac{1}{(n-2m-1)\sqrt{d}}\cdot\min_{m+1\leq i\leq n}\sum_{l\in\Gamma_{w_{i}}^{n-m-2}}D(w_{l},w_{i})+\frac{1}{\sqrt{d}}\cdot\max_{m+1\leq i\leq n}D(w_{i},w_{G})$ where $d$ is the number of parameters in the global model, $\Gamma$ is the set of $n-m-2$ benign local models having the smallest Euclidean distance to $w_{i}$, and $D$ is the Euclidean distance. $\lambda$ is halved until one of the compromised models is selected, or $\lambda$ is less than a threshold.
Untargeted-Med [7] manipulates the model parameters based on the maximum and minimum so that chosen coordinate-wise median values direct toward an inverse direction. The attack starts with defining $w_{max,j}=\text{max}\{w_{(m+1),j},w_{(m+2),j},...,w_{n,j}\}$, which indicates the maximum, and $w_{min,j}=\text{min}\{w_{(m+1),j},w_{(m+2),j},...,w_{n,j}\}$ which indicates the minimum of the $j$th local model parameters on the benign clients, respectively. Also, to avoid sampled $m$ numbers being outliers, if $s_{j}=-1$, $m$ numbers in [$w_{max,j},b\cdot w_{max,j}$] (when $w_{max,j}>0$) or [$w_{max,j},b/w_{max,j}$] (when $w_{max,j}\leq 0$) are randomly sampled. Otherwise, $m$ numbers [$w_{min,j}/b,w_{min,j}$] (when $w_{min,j}>0$) or [$b\cdot w_{min,j},w_{min,j}$] (when $w_{min,j}\leq 0$) are randomly sampled. We set $b=2$ as the same as the paper [7].
3.3.2 Targeted Backdoor Attacks
are based on [3]. Each malicious client owns one sample of mislabeled images and trains the local model on it. We trained the model on backdoor tasks along with the main task(s); the training went on for both malicious and benign tasks to maintain benign accuracy such that backdoor training remains stealthy. Then we boosted malicious clients’ updates to negate the combined effect of the benign agent: $w_{i}^{t}=w_{G}^{t-1}+\alpha_{m}(w_{i}^{t-1}-w_{G}^{t-1})\text{ for }i=1,...,m$ where $t$ is the current epoch and $\alpha_{m}$ is a boosting factor. For DBA [33], we implant split backdoor triggers into the input so that the global model can be attacked in a distributed manner. For more detail, refer to [33]. In CIFAR10, for example, we manipulate the model parameters to misclassify an image of ‘airplanes’ as ‘bird’ whereas normally classifying the image of other classes.
4 FedCC: Robust Aggregation against Poisoning Attacks
4.1 Overview of FedCC
Figure 1 depicts an overview of FedCC. (1) A server initializes and advertises the global model to $n$ participating clients, selected out of total $k$ clients with a probability of fraction $C$. (2) The clients individually train their local models using their respective datasets and send the trained weights to the server for aggregation. Our proposed method intervenes at this moment. (3) The server extracts PLRs from both the global model and each local model, and (4) compares RBF Kernel CKA values between them. (5) The CKA values are then clustered into two, potentially representing benign and malicious clients. (6) The CKA similarity values are then normalized across all clients for every layer as $a$, but the second last layer’s CKA similarity values are normalized within the larger cluster as $b$, backed by the assumption that the number of malicious clients does not exceed $n/2-1$. (7) Finally, the server aggregates each layer of local models weighted by $a$, but the PLR weighted by $b$, to update a global model for the current epoch and proceeds to the next epoch by distributing the updated global model to the newly selected $n$ clients.
Figure 1: An Overview of FedCC. $k$, $n$, $a$, and $b$ indicate the kth layer, the number of participating clients, normalized similarity across all layers (i.e., across_cka), and normalized similarity within the cluster (i.e., within_cka), respectively.
4.2 Detailed Design
A major challenge when filtering malicious clients in the presence of non-IID data is the server’s lack of knowledge regarding the data distribution. Even if the data is benign, local learning models can exhibit significant angular or magnitude differences if they are non-IID. Thus, we require a similarity metric independent of data distribution and not influenced by the distance or direction of model parameters.
4.2.1 Penultimate Layer
is the output of the second-to-last layer before the softmax layer in CNN. In federated learning, where the server cannot access local training data, it receives trained model parameters (weights) from selected clients. Therefore, the collected weights are the only available information for determining whether a client is malicious. PLR takes the end result of the convolution and pooling process that will reach a classification decision. Thus, we assumed that computed weights of the second last layer might contain the essential information that strongly influences classification results.
Figure 2: Cluster Distance of Each Layer
To justify the utilization of PLRs, we measure the distance between two clusters of clients (benign and malicious) using dendrograms with a single linkage metric and correlation distance method. Figure 2 presents cluster distances of each layer under different attack scenarios, such as untargeted-mKrum attacks in IID or non-IID settings. Observably, the cluster distance is the most considerable in PLR to other layers, indicating the highest degree of differences. Based on this observation, we conclude that PLR contains the most distinguishable and indicative information. Therefore, we utilize the weights of the penultimate layer to compare their similarity to the global model.
4.2.2 Kernel CKA[13]
is a similarity metric to assess the similarity or dissimilarity between two neural networks. It compares representations at different layers or models trained in different ways. The sophisticated similarity comparison ability of Kernel CKA makes it a suitable choice for our purpose.
To justify the effectiveness of Kernel CKA, we compared its performance with other similarity metrics such as linear CKA, MMD (not normalized CKA), cosine similarity (angle), and Euclidean similarity (distance). We evaluated the test accuracy of FedCC using the CIFAR10 dataset under two untargeted attacks and one targeted attack, separately in IID and non-IID settings. The results are summarized in Table 1. In the ’targeted’ column, the values represent the combination of test accuracy and backdoor confidence, with higher test accuracy and lower backdoor confidence indicating better performance. We observed that Kernel CKA consistently yields the highest performance across experiments; therefore, we decided to utilize Kernel CKA to compare the similarities of PLRs.
Table 1: Comparison of Performance with Various Similarity Metrics
Method Fang-Med Fang-mKrum Targeted
IID NIID IID NIID IID NIID
Kernel CKA 69.20 41.00 70.22 43.24 71.44/6e-07 54.62/0.0118
Linear CKA 10.00 13.13 64.09 39.55 71.02/0.0007 49.53/0.0616
MMD 63.39 40.90 69.69 32.27 70.85/1e-09 50.51/9e-05
Cosine 68.82 33.90 68.81 10.04 69.76/0.0002 53.66/0.0529
Euclidean 69.06 27.82 68.54 41.57 69.17/0.0221 52.20/0.0015
4.2.3 FedCC
is an aggregation method that combines local clients’ model parameters, based on the Kernel CKA similarities between the global and local models’ PLRs. An overall algorithm is described in Appendix 0.B. A server first extracts PLRs of the global and local models and calculates the kernel CKA between them. We specifically focus on comparing PLRs because they exhibit the most distinguishable differences among differently trained models, as demonstrated in previous sections. Additionally, we utilize RBF Kernel CKA, as it effectively captures similarity differences between non-IID and malicious clients. The similarity values obtained are then clustered using a simple K-means clustering algorithm. Any clustering method that produces only two clusters can be used. Considering the assumption that the number of attackers is less than half of the total number of clients ($n/2$), we designate the clients in the larger cluster as ‘candidates.’
Since the earlier layers in CNNs learn global features and the latter layers embed local features [15, 25, 23], we used two differently normalized CKA similarity values: one normalized within the candidates (i.e., within_cka, denoted as $b$ in Figure 1) and one normalized across all the clients (i.e., across_cka, denoted as $a$ in Figure 1). For the PLR (second-last layer), we apply within_cka to the clients belonging to the larger cluster, whereas for all other layers, we apply across_cka to weighted average model parameters. By doing so, we can ensure that the end magnitude of the aggregated parameters stays similar to before the manipulations, leveraging the fact that the sum of the normalized values is one. Moreover, since our method evaluates clients in every global round and does not exclude any local models, it is also effective for adaptive attacks like DBA.
5 Experiments
5.1 Dataset and Model Architecture
We use three benchmark vision datasets: Fashion-MNIST (fMNIST) [32], CIFAR10 and CIFAR100 [14]. Details about datasets can be found in Appendix 0.A. Considering that the end devices are normally incapable of handling heavy computation due to resource or communication constraints, we used lightweight CNN architectures for the experiments. For fMNIST, we used a 4-layer-CNN with two convolutions and two fully connected (FC) layers, incorporating dropout with probabilities of 0.25 and 0.5 before and after the FC layer, respectively. For CIFAR10, we utilized a 5-layer-CNN with three convolutions, each followed by a max-pool layer and two FC layers, applying dropout with a probability of 0.5 before the first FC layer. For CIFAR100, we adopted LeNet [16].
5.2 Non-IID Simulation
Non-IID is a more feasible assumption considering the diverse and massive nature of data in practice. To standardize non-IID settings, we use Dirichlet distribution, which is widely used and well-known for its capability of representing real-world data [18]. A concentration parameter $\alpha$ controls the degree of non-IID; the smaller $\alpha$ is, the more likely the clients hold examples from only one randomly chosen class. We set the $\alpha$ to 0.2 to represent non-IID, based on [11, 38].
5.3 Experimental Setup and Baselines
We implement two untargeted attacks from [7] and two targeted attacks from [3] and [33], along with defense baselines including Krum, multi-Krum [4], Coomed [34], Bulyan [10], FLTrust [6], FLARE [30], and FedCC. We use ten clients with a 1.0 client participation fraction unless stated otherwise. As in [30], we have three adversaries for each untargeted attack and one for the targeted attack. In the DBA, we randomly select four malicious clients and implant five distributed triggers per batch. To compensate for delayed attack effects, we use FedAvg until epoch 30, after which each defense method is applied. Note that, for a fair comparison, FLTrust is computed using its previous global model without auxiliary data. Benign clients undergo three local epochs, while compromised clients undergo six, with training done using the Adam optimizer (learning rate = 0.001).
5.4 Evaluation Metrics
We evaluated our defense mechanism using two metrics: backdoor confidence (confidence) and global model test accuracy (accuracy). The confidence metric measures misclassification likelihood, ranging from 0 to 1 (lower values are better). Accuracy reflects the global model’s overall test performance under defense methods during attacks, ranging from 0 to 100 (higher values are better). Both metrics were reported under backdoor attacks, with a focus on accuracy for untargeted attacks, as the attacker’s goal is to reduce overall accuracy.
6 Results and Discussion
In this section, we report the performance of our FedCC against two untargeted and two targeted attacks in both IID and non-IID environments. We additionally demonstrate the defense effect against a distributed adaptive attack, DBA [33]. Unlike centralized backdoor attacks, the DBA splits a trigger pattern across multiple adversaries, evading the detection by anomaly detection. We show the result of CIFAR10, with various defense methods applied.
6.1 Non-IID Data Environment
Non-IID is a more feasible assumption considering the diverse and massive nature of data in practice. To standardize non-IID settings, we employ Dirichlet distribution that 0 indicates the most heterogeneity, and 100 mimics homogeneity [20]. A concentration parameter $\alpha$ controls the degree of non-IID; the smaller $\alpha$ is, the more likely the clients hold examples from only one randomly chosen class. Since $\alpha$ being 0.2 represents a highly non-IID scenario based on [11, 38], we set it as such, following [38].
Table 2: Test Accuracy under untargeted attacks in Non-IID setting.
Case data FedAvg Krum MKrum Coomed Bulyan FLTrust FLARE FedCC
Fang -Krum non-IID fM 64.63 16.51 45.91 57.12 13.39 49.37 49.54 66.27
C10 33.69 15.38 20.09 35.70 19.23 36.63 17.03 37.40
C100 2.27 1.00 5.34 4.85 0.98 11.62 7.46 14.51
Fang -Med non-IID fM 16.32 49.33 66.30 68.90 64.12 12.38 52.25 72.76
C10 10.02 25.06 40.52 40.23 32.47 12.39 14.59 42.37
C100 1.00 6.24 13.97 10.27 6.91 1.00 1.00 16.12
6.1.1 Untargeted Model Poisoning Attacks in non-IID setting
Table 2 provides the global model’s accuracy across three datasets under two untargeted (Fang-Krum and Fang-Med) in a non-IID environment. A notable finding is that robust aggregation algorithms often yield lower accuracy than simple averaging (fedAvg). This discrepancy arises from the imperfect identification of malicious clients, leading to the erroneous aggregation of their weights along with those of benign clients. Under Fang-Krum, Krum, Bulyan, and multi-Krum experience significant accuracy drops (16.51%, 13.39%, and 45.91%, respectively, compared to FedAvg’s 64.63%), reflecting the attack’s design to degrade Krum-based defenses. Coomed also struggles, as median weights fail to represent benign clients effectively in non-IID settings. FLTrust mitigates Fang-Krum but performs better on CIFAR10 than fMNIST, likely due to fMNIST’s simpler gradients and lower variance, which challenge its ability to distinguish malicious updates. FLARE’s lower accuracy stems from its approach of scaling the entire model weights uniformly, which ignores knowledge contributed by individual clients. In contrast, FedCC, which selectively averages weights layer-wise, achieves superior performance by boosting or minimizing weights based on their alignment with benign updates. FedCC demonstrates the highest accuracy (66.27%, 37.40%, and 14.51% for fMNIST, CIFAR10, and CIFAR100, respectively) and shows significant improvement in CIFAR100 experiments.
Turning to the Fang-Med attack, FedAvg experiences a significant drop in accuracy due to the attack’s creation of outliers by deviating from median values. Consequently, Coomed and Bulyan perform relatively well since their coordinate-wise median and trimmed mean methods disregard outliers effectively. FLARE continues to perform poorly for a similar reason mentioned earlier. Between the Krum-based methods, multi-Krum and Coomed are more robust than Krum, as they aggregate multiple local models, providing better resilience to outliers. Similarly, Bulyan’s averaging of multiple clients enhances robustness but falls short of Coomed, particularly in non-IID environments. Finally, FedCC achieves the highest accuracy (72.76%, 42.37%, and 16.12% for each dataset) without sharing raw data, showcasing its exceptional performance. This success is attributed to two factors: (1) highly distinguishable information within PLRs and (2) higher CKA similarity between benign clients than between benign and malicious clients.
6.1.2 Targeted Backdoor Attacks
Table 3: Test Accuracy under targeted attacks in Non-IID setting.
Case data FedAvg Krum MKrum Coomed Bulyan FLTrust FLARE FedCC
Target non-IID fM 75.65 45.27 65.97 71.70 57.96 61.82 64.31 75.66
C10 36.16 14.98 30.72 48.97 40.11 44.06 10.18 51.56
C100 4.46 6.18 6.90 12.04 11.16 12.95 1.14 15.26
DBA C10 38.56 24.94 7.09 44.45 34.19 51.49 38.73 52.28
Figure 3: Confidence of Backdoor Task for targeted attacks in Non-IID settings.
Table 3 summarizes test accuracy under targeted attacks in a non-IID setting. To further illustrate the impact, Figure 3 provides a visual representation of backdoor confidence. It is evident that Krum, which selects a single client’s weights as the global model, shows reduced robustness compared to methods that aggregate weights from multiple clients. In contrast, Coomed, multi-Krum, and Bulyan exhibit higher test accuracy. FLARE reduces main task accuracy for reasons similar to those observed in previous experiments. FedCC achieves the highest test accuracy. Since targeted attacks aim to misclassify a specific target class while maintaining the main task accuracy, the best performance of FedCC is attributed to leveraging knowledge from earlier stages of training. This result highlights the superiority of using sophisticated weighted knowledge and filtering malicious clients via CKA similarity, which outperforms first-order statistical methods like mean or median. A similar trend is observed with the DBA attack, except for multi-Krum; the lowest accuracy occurs because all four malicious clients, each with a distributed trigger, are treated as outliers and excluded from aggregation.
Notably, Figure 3 demonstrates that FedCC significantly reduces backdoor confidence. Note that since the backdoor confidence of DBA fluctuates a lot due to its distributed nature, we omit it for brevity. Unlike other methods, such as FLTrust or multi-Krum, which exhibit fluctuating confidence values, FedCC maintains consistently low confidence, emphasizing its resilience and independence from client-specific data distributions. This advantage stems from the effective utilization of CKA, enabling similarity calculations even when models are trained on different datasets with varying distributions. While other baselines, like Bulyan or FLTrust, also reduce confidence, we argue that FedCC not only preserves but potentially increases accuracy by precisely distinguishing malicious clients.
6.2 IID Data Environment
In an IID setting, we evenly divide the training dataset among all clients so that each client’s data distribution is identical and of the same size.
Table 4: Test Accuracy under untargeted attacks in IID setting.
Case data FedAvg Krum MKrum Coomed Bulyan FLTrust FLARE FedCC
Fang -Krum IID fM 75.55 31.66 87.78 87.62 50.30 89.53 79.16 89.57
C10 49.67 40.86 63.42 57.40 12.67 68.25 25.77 69.84
C100 13.72 1.04 7.64 6.17 1.59 17.14 7.49 18.47
Fang -Med IID fM 20.86 85.33 89.53 86.70 87.45 21.36 71.08 89.66
C10 9.51 54.28 69.68 59.20 57.69 9.92 49.82 70.52
C100 0.87 12.27 16.52 14.43 12.56 1.16 5.83 17.83
6.2.1 Untargeted Model Poisoning Attacks
Table 4 shows the test accuracy of the global model trained on three datasets under untargeted attacks in IID settings. While trends in these experiments are similar to those in the non-IID setting, the accuracy values are slightly higher due to the more consistent data distribution across and updates from clients in the IID case. A notable deviation from the non-IID experiments is when Coomed is applied: the accuracy increases from 75.55% (FedAvg) to 87.62% (Coomed). This indicates that the coordinate-wise median values closely align with the global model’s weights with minimal deviation and represent benign models well, as the adversary’s impact is limited due to the majority of benign clients. FedCC achieves the highest accuracy across all methods (89.57%, 69.84%, and 18.47% for each dataset) due to its effective CKA similarity measurement and sophisticated layer-wise aggregation strategy.
Under the Fang-Med attack, the accuracy of FedAvg experiences a significant drop (89.89% to 20.86%) due to the peculiarity of outliers when the clients have IID data. Similar to the result of Fang-Krum, Krum, Coomed, and Bulyan demonstrate similar accuracy values, indicating that coordinate-wise median values indeed represent benign clients’ parameters. FLTrust, however, is less effective in mitigating the Fang-Med attack than Fang-Krum, even in the IID setting; it then means that the performance difference is not due to data distribution but rather its inherent vulnerability. In Fang-Krum, malicious updates are deliberately far from the global model, allowing FLTrust to identify discrepancies easily. In contrast, the malicious updates in Fang-Med are subtler and only slightly deviate from the global model or median, making it harder for FLTrust to distinguish them from benign updates. Meanwhile, FedCC consistently achieves the highest accuracy (89.66%, 70.52%, and 17.83% for each dataset), demonstrating its capability to mitigate attacks involving more nuanced changes to model updates without extreme outliers.
6.2.2 Targeted Backdoor Attacks
Table 5: Test Accuracy under targeted attacks in IID setting.
Case data FedAvg Krum MKrum Coomed Bulyan FLTrust FLARE FedCC
Target IID fM 88.27 86.63 87.03 89.41 89.45 89.59 75.29 90.01
C10 64.68 57.69 71.19 69.85 68.76 68.61 11.09 71.64
C100 13.83 5.72 17.57 13.24 15.11 17.08 1.03 18.61
DBA C10 10.00 35.58 51.40 10.00 16.16 56.69 10.00 58.04
Figure 4: Confidence of Backdoor Task for targeted attacks in IID settings.
Table 5 and Figure 4 present the test accuracy and backdoor confidence under targeted backdoor attacks in IID settings. Similar to the untargeted attack scenario, the general test accuracy in IID settings is higher than in non-IID settings.
Reducing backdoor confidence is as critical as maintaining main task accuracy, as attackers aim to stealthily embed backdoor tasks. A key observation is FedCC’s remarkable ability to reduce backdoor confidence to near zero while preserving high main task accuracy. In contrast, the fluctuating backdoor confidence observed in Multi-Krum and Flare is largely due to the distributed trigger, which confuses their defense mechanisms and hampers their ability to filter out outliers effectively. Other methods, such as Multi-Krum, Flare, and FLTrust, show significant variability in backdoor confidence, further highlighting FedCC’s superiority.
Notably, FedCC achieved the highest main task accuracy across both targeted backdoor attacks and distributed backdoor attacks (DBA), consistently neutralizing the backdoor task. These results underscore FedCC’s ability to mitigate targeted backdoor attacks through precise and efficient integration of prior knowledge to preserve main task performance while zeroing out the attack confidence.
6.3 Performance in Varying FL Settings
We additionally measured the effectiveness of FedCC in various non-IID settings, including different numbers of malicious clients, fractions of participation, and numbers of local epochs. Since Coomed was as effective as ours, and FLARE also used PLRs, we compared three methods to FedCC, including FedAvg.
6.3.1 Impact of the Number of Malicious Clients
Figure 5: Test Accuracy with different numbers of malicious clients
We assessed the impact of different numbers of malicious clients under untargeted attacks, specifically Untargeted-Krum and Untargeted-Med. In this set of experiments, we kept the total number of clients and participation fraction fixed at 10 and 1, respectively. To adhere to the assumption that the number of attackers is less than half of the participating clients, we investigated the test accuracy considering scenarios involving up to four malicious clients.
Figure 5 illustrates the accuracy in the given experimental settings, and it is evident that FedCC outperforms the other defense methods. We observe a general trend where the accuracy of defense methods decreases as the number of attackers increases. Notably, Coomed experiences the most significant drop in accuracy when transitioning from one malicious client to four malicious clients. This can be attributed to the fact that as the participation of malicious clients increases, there is a higher likelihood of the median values being influenced by their malicious contributions.
In contrast, FedCC does not rely solely on geometric measures such as angles or distances but instead leverages hidden correlations between networks to identify and filter out malicious clients. As a result, FedCC demonstrates superior accuracy in filtering out malicious clients, regardless of the number of attackers involved. It is important to note that the accuracy degradation observed as the number of malicious clients increases primarily due to aggregating fewer clients rather than the malicious clients themselves being selected. These findings underscore the robustness and effectiveness of FedCC in defending against untargeted attacks, as it consistently outperforms other defenses across varying numbers of malicious clients.
6.3.2 Impact of Fraction
Table 6: Test Accuracy under untargeted attacks with different fractions of clients being selected.
Untargeted-Krum Untargeted-Med
Frac Data FedAvg Med FLARE FedCC FedAvg Med FLARE FedCC
0.1 fM 55.31 49.83 34.02 64.83 16.57 66.41 52.24 69.52
C10 10.06 22.55 14.50 20.49 10.00 15.33 10.00 29.81
0.3 fM 64.22 57.52 10.00 73.55 16.26 58.07 10.00 61.12
C10 24.24 12.59 10.00 27.81 10.98 22.61 10.00 38.27
0.5 fM 62.37 58.20 10.00 76.41 18.36 62.49 10.00 69.05
C10 23.99 17.28 10.06 34.32 9.87 27.83 10.00 37.27
We examine the impact of the participating fraction of clients where the number of total clients is fixed at 100. The fractions are 0.1, 0.3, and 0.5, such that the numbers of clients being aggregated are 10, 30, and 50, respectively. The proportion of malicious clients is fixed at 0.3, such that the malicious clients are 3, 9, and 15, respectively. Table 6 summarizes the accuracy. We can observe that the accuracy tends to increase as the participation fraction is greater under untargeted-Krum attacks. It implies that the more client participation in a non-IID setting, the more accurate the global model is if the attacks are mitigated. Under untargeted-Med attacks, the accuracy is the highest with FedCC and second highest with Coomed. Throughout the experiment, FLARE prevented the global model from convergence, and the test accuracy fluctuated. The fluctuation is presumably due to the weight scaling while randomly selecting clients. Different from previous experiments, the server chooses a fraction of the client every round, meaning not-yet-trained models can be selected. In this environment, weight scaling leveraged by FLARE constantly improperly scales the local weights, causing divergence and fluctuation. These experimental results indicate that selectively choosing clients based on CKA similarity is more reliable than taking the median value of each coordinate in a non-IID setting. We also observed that as the participation fraction grows, the accuracy grows. This is due to the rising number of aggregated local models; the more models are aggregated, the more general global model is generated.
7 Conclusion and Future Work
FL has emerged in response to the rising concerns about privacy breaches while employing AI techniques. FL, however, is vulnerable to model poisoning attacks due to a server’s blindness to local datasets, making it difficult to assume data distribution and model parameter integrity. In line with this, many robust aggregation algorithms and defenses have been proposed; still, existing approaches are orthogonal to various attacks and lack consideration of non-IIDness. Throughout the exhaustive experiments, FedCC mitigates both untargeted and targeted (or backdoor) attacks while demonstrating its effectiveness in non-IID data environments. We leave theoretical guarantees of FedCC and experiments when clients have different neural network architectures or tasks as future work.
Appendix
Appendix 0.A Dataset and Model Architecture
We use three benchmark vision datasets: Fashion-MNIST (fMNIST) [32], CIFAR10, and CIFAR100 [14].
fMNIST consists of 60,000 training and 10,000 testing samples of 28$\times$28 grayscale images and associating labels of 10 classes. The labels include T-shirt/top, Trouser, Pullover, Dress, Coat, Sandal, Shirt, Sneaker, Bag, and Ankle boot.
CIFAR10 consists of 50,000 training and 10,000 testing examples of 32$\times$32 color images and associating labels of 10 classes. The classes include airplane, automobile, bird, cat, deer, dog, frog, horse, ship, and truck.
CIFAR100 is the same as CIFAR10 except for the categorization. It is labeled as 100 classes, each containing 600 images. Thus, there are 500 training and 100 testing images per class. Each class (i.e., ‘fine’ label) belongs to one of the 20 superclasses.
Appendix 0.B FedCC: Algorithm
Algorithm 1 summarizes FedCC’s workflow.
Algorithm 1 FedCC
Input: global_w, n local_w ($w_{1}$, …, $w_{n}$)
      Output: agg_w, larger_cluster_members
$\triangleright$ Get PLRs of local models
1:for $i<n$ do
2:     local_plrs[$i$] $\leftarrow$ local_w[$i$][second_last_layer]
3:end $\triangleright$ Extract global PLR
4:glob_plr $\leftarrow$ global_w[second_last_layer] $\triangleright$ Apply kernel CKA to the plrs
5:cka[$i$] $\leftarrow$ kernel_CKA(glob_plr, local_plrs[$i$]) $\triangleright$ Normalize similarities globally
6:across_cka $\leftarrow$ normalize(cka) $\triangleright$ Apply Kmeans clustering algorithm
7:kmeans $\leftarrow$ kmeans(n_clusters=2, cka)
8:labels $\leftarrow$ kmeans.labels_
9:count $\leftarrow$ counter(labels)
10:larger_cluster $\leftarrow$ 1
11:if count[00] $>$ count[$1$then
12:     larger_cluster $\leftarrow$ 0
13:end $\triangleright$ Identify larger cluster members
14:larger_cluster_members = where(labels == larger_cluster) $\triangleright$ Normalize similarities within larger cluster
15:within_cka $\leftarrow$ normalize(cka[larger_cluster_members])
16:Initialize agg_weights as zero tensor $\triangleright$ Differently weigh weights per layer
17:for layer in model_layers do
18:     if layer is up to second-to-last layer then
19:         agg_weights[layer] $\leftarrow$ weighted average using across_cka
20:     else if layer is second-to-last layer then
21:         agg_weights[layer] $\leftarrow$ weighted average using within_cka
22:     else
23:         agg_weights[layer] $\leftarrow$ weighted average using across_cka
24:     end if
25:end for
26:return agg_weights, larger_cluster_members